Write a quadratic equation in standard form that has the solutions 1 and -8
Answer by ewatrrr(24785)

 
Hi,
Write a quadratic equation in standard form that has the solutions 1 and -8
y = (x-1)(x+8) = x^2 -7x - 8
y = x^2-7x-8 |y = ax^2+bx+c